OpenStudy (anonymous):

Describe geometrically the regions of the complex plane given by each of the following sets of complex numbers:

OpenStudy (anonymous):

\[\frac{ \pi }{ 2 }<\arg(\frac{ z-1 }{ z })\le \pi\]

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@smokeydabear

OpenStudy (nathan917):

@SerikMB Your question as stated doesn't really make any senses since z - 1z = z - z = 0. I'm going to guess that you actually meant z - iz. Then I'm not sure what you mean by π2 so I'm going to guess that you mean π times 2 which I would usually write as 2π z - iz = z(1 - i) Write z in polar form z = r e^(iθ) write (1 - i) in polar form (1 - i) = √2 e^(-iπ/4) hence z - iz = r√2 e^i(θ - π/4) The argument is θ - π/4 Then the requirement is: 2π < θ - π/4 < π (2+1/4)π < θ < (1+1/4)π 9π/4 < θ < 5π/4 I asked it on yahoo answer. 4u

OpenStudy (nathan917):

Got what you wanted?